OpenStudy (anonymous):

Write the equation of each parabola described. 1. Vertex at (2,5). Focus at (2,-3) 2. Directix:x=3/4. Focus at (5/4,-3)

OpenStudy (anonymous):

Vertex ( h, k) = ( 2, 5 ) and F (2, -3) -> ( x - h )² = 4p ( y - k) -> ( x - 2 )² = -32 ( y - 5 )

OpenStudy (anonymous):

2. Directrix :x = 3/4. Focus at (5/4,-3) ( y - k)² = 4p ( x - h ) ( y + 3)² = ( x - 1)